Sol Summary – January 9th

Crew 290 Sol Summary Report 09-01-2024

Sol: 2

Summary Title: Science not working out, but we are!

Author’s name: Madelyn Hoying

Mission Status: Nominal

Sol Activity Summary:

Today kicked off our schedule for exercise countermeasures, with the crew taking 1.5-hour blocks each day to complete a HIIT- and resistance-based workout plan. The first EVA kicked off smoothly with Rebecca, Ben, and Nicole completing the Marble Ritual excursion while Madelyn was CAPCOM, Anna evaluated an issue with Suit #8’s airflow, and Anja prepared REMI for operation. A slight contingency with Becka’s headset falling down on EVA 02 resulted in a much-preferred headset placement for Madelyn on EVA 03, resolving the comms difficulties faced in EVA 01. After a lunch of alfredo pasta prepared by Ben, EVA 03 served as a systems check of REMI to enable data collection in later excursions. Becka and Anna spent the afternoon troubleshooting the pH meter and other Science Dome equipment for Project MADMEN. Anja prepped a dinner of taco bowls (buddha-FULL buddha bowls, according to Anna) and the crew is preparing for a game night tonight, with a possible call sign ceremony. Stay tuned to see our call signs tomorrow!

Look Ahead Plan: pH meter solution, EVA 04 for REMI data collection and to begin sample collection for MADMEN

Anomalies in work: Some science hiccups along the way with pH meter functionality not established, and a solvable issue in REMI operations

Weather: Cloudy compared to yesterday’s clear skies, but warm enough for just two layers with the flight suits by EVA 03

Crew Physical Status: nominal

EVA: 02, Marble ritual completed with Becka, Ben, and Nicole; 03, REMI testing with Madelyn, and Anja

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, EVA Reports, EVA Request

Support Requested:

1. Are there any additional pH meters, or parts to fix Shannon’s? After troubleshooting, the crew is not able to get hers operational.

2. Crew changed the batteries on Shannon’s pH meter – what is the procedure for disposal of the 3V coin battery that was removed?

3. Wiring diagrams for the one-piece suits are requested, to check for a solution to suit #8’s lack of airflow (see ops report for more detail).

4. Would it be possible to build up a bit of dirt around the ramp out of the RAM airlock, so that the crew does not have to lift REMI over the lip each time? This could be done as an EVA task if needed.

Copyright © The Mars Society. All rights reserved. | Main Site